# DeCo-SGD: Joint Optimization of Delay Staleness and Gradient Compression Ratio for Distributed SGD

### v1

@@ B.3 Proof of Theorem 2
$X_{t+1}$ $=X_{t}+\gamma^{2}\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\rVert^{2}-\underbrace{2\gamma\mathbb{E}\left\langle\hat{\mathbf{x}}_{t}-x^{*},\mathbf{g}_{t}\right\rangle}_{:=A_{2}}.$
@@
The second line is due to the reason that $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L$-smooth and $\mu$-strongly convex. Then choosing $\gamma\leq\frac{1}{4L},\mu\leq L$, we have
$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}+(2L\gamma^{2}-\gamma)F_{t}+\gamma(4L+2\mu)(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2})$
$\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}f_{t}+6L\gamma(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
@@
Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $\tilde{B}_{t}$ as in Eq. 3.2. Then with $\gamma\leq\frac{1}{8\sqrt{L\tau}}$, we have
$\sum_{t=0}^{T}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n},$
where $W_{t}=\sum_{i=1}^{n}w_{i}$.
Proof. This lemma follows from the Lemma 20 in the previous work [14].
@@
Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $B_{t}$ as in Eq. 3.2. Then by using the relative compressor, with $\gamma\leq\frac{1}{16\sqrt{L\phi}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$ we have
$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert B_{t}\rVert^{2}$ $\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2}),$
where $W_{t}=\sum_{i=1}^{n}w_{i}$.
proof. Denote $E_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$, and we have:
$w_{t}E_{t}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}w_{i}(1+\frac{\delta}{4})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2})$
$+\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{i=0}^{t}(1-\frac{\delta}{4})^{t-i}w_{i}\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}]$
$\sum_{t=0}^{T}w_{t}E_{t}$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{j=0}^{\infty}(1-\frac{\delta}{4})^{j}\sum_{i=0}^{T}w_{i}\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})]$
$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}[\frac{4}{\delta}\sum_{i=0}^{T}w_{i}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})].$
Choosing $\frac{4\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we complete the proof.
@@
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex. Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{8L\sqrt{\tau}},\frac{1}{16L\sqrt{\phi}}\}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L}{\mu}+\frac{\sqrt{L\tau}}{\mu}+\frac{\sqrt{L\phi}}{\mu})\end{split}$
iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\min\left\{\frac{\mu\gamma}{2},\frac{\delta}{4}\right\})^{-t}$.
@@
B.4 Proof of Theorem 3

### v2

@@ B.3 Proof of Theorem 2
$X_{t+1}$ $=X_{t}+\gamma^{2}\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\rVert^{2}-\underbrace{2\gamma\mathbb{E}\left\langle\hat{\mathbf{x}}_{t}-\mathbf{x}^{*},\mathbf{g}_{t}\right\rangle}_{:=A_{2}}.$
@@
The second line is due to the fact that $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L$-smooth and $\mu$-strongly convex. Then choosing $\gamma\leq\frac{1}{4L},\mu\leq L$, we have
$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}+(2L\gamma^{2}-\gamma)F_{t}+\gamma(4L+2\mu)(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2})+\gamma^{2}\sigma^{2}$
$\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}F_{t}+6L\gamma(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2})+\gamma^{2}\sigma^{2}.$
@@
Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $\tilde{B}_{t}$ as in Eq. 6. Then with $\gamma\leq\frac{1}{4\sqrt{2}\ L\tau}$, we have
$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}F_{t}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n},$
where $W_{t}=\sum_{i=0}^{t}w_{i}$.
Proof. Based on Eq. 18, when $\gamma\leq\frac{1}{4\sqrt{2}L\tau}$, we have
$\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$ $\leq\frac{1}{128L^{2}\tau}\sum_{j=1}^{\tau}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-j})\rVert^{2}+\frac{\tau\gamma^{2}\sigma^{2}}{n}$
$\leq\frac{1}{64L\tau}\sum_{j=1}^{\tau}F_{t-j}+\frac{\tau\gamma^{2}\sigma^{2}}{n}.$
The second inequality is due to the assumption of L-smoothness $\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert\leq 2LF_{t}$.
We observe that when $\tau\geq 1$, we have $w_{t}=(1-\frac{\mu\gamma}{2})^{-(t+1)}$, $1-\frac{\mu\gamma}{2}\geq 1-\frac{\mu}{8\sqrt{2}\ L\tau}\geq 1-\frac{1}{8\sqrt{2}\ \tau}\leq(1+\frac{1}{2\tau})^{-1}$, thus having $w_{t}=w_{t-j}(1-\frac{\mu\gamma}{2})^{-j}\leq w_{t-j}(1+\frac{1}{2\tau})^{j}\leq w_{t-j}(1+\frac{1}{2\tau})^{\tau}\leq w_{t-j}\exp[1/2]\leq 2w_{t-j}$. Based on this inequality, we have
$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$ $\leq\frac{1}{32L}\sum_{t=0}^{T}\frac{w_{t}}{2\tau}\sum_{j=1}^{\tau}F_{t-j}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n}$
$\leq\frac{1}{32L}\sum_{t=0}^{T}\frac{1}{\tau}\sum_{j=1}^{\tau}w_{t-j}F_{t-j}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n}$
$\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}F_{t}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n}.$
@@
Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $B_{t}$ as in Eq. 6. Then by using the relative compressor, with $\gamma\leq\frac{1}{16LZ\sqrt{2\phi/\delta}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$ we have
$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert B_{t}\rVert^{2}$ $\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}F_{t}+2\gamma^{2}\phi W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2}),$
where $W_{t}=\sum_{i=0}^{t}w_{i}$.
proof. We denote $E_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$, and based on Eq. 19, then we have:
$w_{t}E_{t}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}w_{i}(1+\frac{\delta}{4})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2})$
$\quad+\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{i=0}^{t}(1-\frac{\delta}{4})^{t-i}w_{i}\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}],$
$\sum_{t=0}^{T}w_{t}E_{t}$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{j=0}^{\infty}(1-\frac{\delta}{4})^{j}\sum_{i=0}^{T}w_{i}\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})]$
$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}[\frac{4Z^{2}}{\delta}\sum_{i=0}^{T}w_{i}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})]$
$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}[\frac{8LZ^{2}}{\delta}\sum_{i=0}^{T}w_{i}F_{i}+W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})].$
Choosing $\frac{16LZ^{2}\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{32L}$, we complete the proof.
@@
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex. Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4\sqrt{2}\ L\tau},\frac{1}{16LZ\sqrt{2\phi/\delta}}\}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L(\tau+Z\sqrt{\phi/\delta})}{\mu})\end{split}$
iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\frac{\mu\gamma}{2})^{-t}$.
@@
B.4 Proof of Convergence rate of D-SGD with delayed aggregation
The update rule of D-SGD with delayed aggregation is: $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i}$. Based on the virtual sequence, we have
$\tilde{B}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{\tau}\mathbf{g}_{t-\tau}^{i}{,}\quad\hat{\mathbf{x}}_{t+1}=\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\mathbf{g}_{t}^{i}.$ (20)
Simlar to Section B.2, we have following lemmas.
Lemma 11.
Let $\left\{\mathbf{x}_{t},\hat{\mathbf{x}}_{t},\tilde{B}_{t}\right\}_{t\geq 0}$ be defined in D-SGD with delayed aggregation (Eq. 20), with gradient oracle $\left\{\mathbf{g}_{t}\right\}_{t\geq 0}$ and let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. If $\gamma\leq\frac{3}{2L}$, then we have
$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\frac{\gamma}{2}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{t}\right)\right\|^{2}+2\gamma L^{2}\mathbb{E}\left\|\tilde{B}_{t}\right\|^{2}+\frac{\gamma^{2}L\sigma^{2}}{2n}.$
Proof. We begin with the definition of $\hat{\mathbf{x}}_{t+1}$ and have
$\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\underbrace{\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle}_{:=A_{1}}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\gamma^{2}\sigma^{2}}{2n}.$
We estimate $A_{1}$ as
$A_{1}$ $=\gamma(\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right)-\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle-\mathbb{E}\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{g}_{t}\rangle)$ (22)
$\leq\frac{c_{1}\gamma L^{2}}{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\frac{\gamma}{2c_{1}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}-\gamma\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}$
$\leq-\frac{7\gamma}{8}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma L^{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}.$
The second line is due to Jensen’s inequality (i.e., Eq. 13). We let $c_{1}=4$ in the third line. We bring $A_{1}$ into Eq. 21 and complete the proof.
Theorem 4 (Non-convex convergence rate of D-SGD with delayed aggregation).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. Then there exists a stepsize $\gamma\leq\frac{1}{4L\tau}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\frac{\tau}{n}\sigma^{2}}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$
iterations of D-SGD with delayed aggregation, it holds $\mathbb{E}\lVert\nabla f(\mathbf{x}_{out})\rVert^{2}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, where $\mathbf{T}$ denotes the total iterations, chosen at random uniformly.
Proof. Let $G_{t}:=\sum_{i=0}^{t}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{i}\right)\right\|^{2}$. According to above lemmas (Lemma 11, 6), we have
$\frac{3\gamma}{8}G_{T}\leq\mathbb{E}f(\mathbf{x}_{0})-f^{*}+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}+\gamma^{3}L^{2}(T+1)\frac{2\tau\sigma^{2}}{n}.$
In this way, we have $\gamma\leq\min\{\frac{3}{2L},\frac{1}{4L\tau}\}\leq\frac{1}{4L\tau}$. According to Lemma 3, we let $B=3/8$, $C=\frac{L\sigma^{2}}{2n}$, $D=\frac{L^{2}\tau\sigma^{2}}{n}$, $E=4L\tau$, $r_{0}=\mathbb{E}f(\mathbf{x}_{0})-f^{*}$, completing the proof.
Lemma 12.
Let $\left\{\mathbf{x}_{t},\hat{\mathbf{x}}_{t},\tilde{B}_{t}\right\}_{t\geq 0}$ be defined in D-SGD with delayed aggregation (Eq. 20), with gradient oracle $\left\{\mathbf{g}_{t}\right\}_{t\geq 0}$ and let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-strongly convex. If $\gamma\leq\frac{1}{2L}$, then we have
$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}F_{t}+\gamma^{2}\sigma^{2}+5\gamma L\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}.$
Proof.
$X_{t+1}$ $=X_{t}+\gamma^{2}\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\rVert^{2}-\underbrace{2\gamma\mathbb{E}\left\langle\hat{\mathbf{x}}_{t}-\mathbf{x}^{*},\mathbf{g}_{t}\right\rangle}_{:=A_{2}}.$
We estimate the term $A_{2}$ as
$A_{2}$ $=-2\gamma(\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{x}_{t}-\mathbf{x}^{*}\right\rangle-\ \mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\right\rangle)$
$\leq-\gamma\mu\mathbb{E}\lVert\mathbf{x}_{t}-\mathbf{x}^{*}\rVert^{2}-2\gamma F_{t}+\frac{1}{2}\gamma F_{t}+4\gamma L\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$
$\leq-\frac{\gamma\mu}{2}X_{t}-\frac{3}{2}\gamma F_{t}+\gamma(4L+\mu)\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}.$
The second line is due to the fact that $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L$-smooth and $\mu$-strongly convex. Then choosing $\gamma\leq\frac{1}{2L},\mu\leq L$, we have
$X_{t+1}\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}F_{t}+5L\gamma\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\gamma^{2}\sigma^{2}.$
This completes the proof.
Theorem 5 (Convex convergence rate of D-SGD with delayed aggregation).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-strongly convex. Then there exists a stepsize $\gamma\leq\frac{1}{4\sqrt{2}\ L\tau}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L\frac{\tau}{n}\sigma^{2}}}{\mu\epsilon^{1/2}}+\frac{L\tau}{\mu})\end{split}$
iterations of D-SGD with delayed aggregation, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\frac{\mu\gamma}{2})^{-t}$.
Proof. By substituting Lemma 9 into Lemma 12, we can get
$X_{T}\leq w_{T}X_{0}-\frac{\gamma}{2}\sum_{i=0}^{T}w_{t}F_{t}+\frac{W_{T}\gamma^{2}\sigma^{2}}{n}+5\gamma L\sum_{i=0}^{T}w_{t}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$
$\leq w_{T}X_{0}-\frac{11\gamma}{32}\sum_{i=0}^{T}w_{t}F_{t}+W_{T}\gamma^{2}\sigma^{2}+5LW_{T}\gamma^{3}\frac{\tau\sigma^{2}}{n}.$
According to Lemma 4, we complete the proof.
B.5 Proof of Theorem 3